ableVision has been approached by the City of Mirada to run its cable operations in 2020. After negotiating with key parties, CableVision has made the following agreements: It will offer Mirada residents a basic set of 25 cable television stations at a rate of $32.99 per month. It will pay the city $110,000 per month plus $3.25 of the monthly cable subscriber rate per month to maintain the physical facilities. It will actually pay another company an annual fixed fee of $730,000 plus $8.00 per cable subscriber per month to broadcast the 25 channels. CableVision estimates that operating costs for billing, program news mailings, etc. will be $110,000 per month plus 7% of the monthly subscriber rate. CableVision has several questions about its monthly revenues, costs, and profits in 2020.


Part A (8 tries; 8 points)
1. What is the estimated monthy contribution margin per cable subscriber for CableVision in 2020?   

2. What are the estimated total monthly fixed costs for CableVision in 2020?   



Part B (8 tries; 8 points)
1. What is CableVision's estimated monthly operating income in 2020 if 20,000 residents subscribe?   

2. How many monthly subscribers would be required for CableVision to break even in 2020?   

3. How many monthly subscribers would be required for CableVision to earn $24,000 per month in 2020?   

4. Assuming a tax rate of 35%, how many monthly subscribers would be required for CableVision to earn $24,000 per month in 2020?   

Part C (4 tries; 4 points)
Some of CableVision's managers are uncertain about their estimate of monthly fixed operating costs. Assuming that 22,000 residents subscribe, how large can monthly fixed operating costs be for CableVision to still earn $24,000 per month in 2020 (ignore taxes)?   
Answer #1

1.Monthy contribution margin per subscriber = Revenue per subscriber – variable costs per subscriber

= 32.99-3.25-8-7%*32.99

= $19.4307

i.e. $19.43

2.Estimated monthly fixed costs = 110,000+730,000/12 + 110,000

= $280,833.33

i.e. $280,833

1.Operating income = Contribution margin – fixed costs

= 20000*19.43 – 280,833

= $107,767

2.Break even subscribers = Fixed costs/Contribution margin per subscriber

= 280,833/19.43

= 14,453.58 subscribers

i.e. 14,454 subscribers

3.Required = (24000+280,833)/19.43

= 15,689 subscribers

4.Required = (24000/0.65 + 280,833)/19.43

= 16,354 subscribers

1.Contribution Margin earned = 19.43*22000 = $427,460

Required profit = 24000

Other fixed costs = 110000 + 730,000/12 = $170,833

Maximum monthly operating fixed costs = $427,460-24000-170,833

= $232,627